# Fermionic tensor network contraction for arbitrary geometries

### v2

@@ II.1 Standard tensor networks and symmetries
We will consider only symmetric tensors with well-defined total irrep. To define how the total irrep is obtained from the irreps of the indices, we associate the indices of the tensor with an Abelian signature, a sequence of pluses and minuses. For example, for $A_{ijkl}$, the ++++ signature means
$A_{ijkl}\neq 0\ \mathrm{if}\ s(i)+s(j)+s(k)+s(l)=s(A)$ (4)
where $s(A)$ is the total irrep of $A$. Any replacement of $+$ by -- in the signature yields the equivalent replacement in Eq. (4). Note that for finite Abelian groups (e.g., $\mathbb{Z}_{2}$), addition and subtraction is modulo $H$.
@@
The unordered contraction in Eq. 2 defines the value of $\mathcal{T}$. However, the cost to compute $\mathcal{T}$ further depends on the order of the tensor contractions. If approximate contraction is used, both cost and accuracy also depend on when and which shared bonds are compressed. The order of tensor contraction in the exact case defines a contraction tree (see Fig. 1(a)) where the vertical axis illustrates the time-ordering of the pairwise contractions. In the approximate contraction case, Fig. 1(b), the contraction tree contains additional lines between branches indicating the compression of bonds between tensors.
@@ II.2 Fermionic tensors and contractions
A fermionic tensor network is formed from fermionic tensors. We thus first discuss the properties of fermionic tensors. In our implementation, these are stored as Abelian arrays of numbers with additional metadata for parity and other symmetries, and the internal index order. Parity is a $\mathbb{Z}_{2}$ symmetry with irreps $\sigma\in\{0,1\}$; for index $i$, we write $p_{i}$ as the parity of $i$. We will assume the fermionic tensors have well-defined total parity 00 or $1$. Fermionic tensors may have additional symmetries, which may or may not be independent of parity. A common symmetry is $\mathbb{U}(1)$ symmetry, associated with fermion number or projected spin number. In this case, odd integer irreps of $\mathbb{U}(1)$ have parity $1$, while even integer irreps of $\mathbb{U}(1)$ have parity 00.
@@ II.4 Globally ordered formalism
where we store the additional data on the initialization order in the form of an integer position for each fermionic tensor (the superscripts in Eq. 14). $\mathcal{T}$ is then obtained by contracting the network in sequence from left to right. Transposition of $\mathcal{T}$ (e.g., when taking the hermitian conjugate to go from $|\Psi\rangle$ to $\langle\Psi|$ involves reversing the order of the tensors appearing in Eq. 14, as well and reversing the indices of all the tensors.
@@
In this DAG picture, we can also visualize what happens when one changes the global ordering. When moving tensors past each other to construct a new global order, one needs to construct a new DAG corresponding to the new order (see sequence in Fig. 3a). In general, this requires the reversal of multiple arrows in the original DAG, and these reversals correspond to the bond parity tensors that appear in the algebraic discussion.
@@ II.5 Locally ordered formalism
From here is a short leap to the locally ordered formalism. It is clear from the above that it is not necessary to limit ourselves to DAGs, but one can use other arrow graphs, where the arrows only define the local ordering of two tensors in a given contraction. In this case, the value of the fermionic tensor network is given by
@@
where $\mathcal{G}$ denotes the fermionic contraction following the additional data where every contracted bond has an arrow, and there is no longer any global ordering of the tensors. The computation of the bond parity is needed only on the bond between two tensors when they are contracted: if they are contracted in the direction of the arrow that bonds them, there is no parity, while if they are contracted in the opposite order, we compute the bond parity (or equivalently, first reverse the arrow). The above assumes that we are working with even parity tensors, so there is no global phase. However, odd parity tensors can be implemented by using an additional dummy (size 1) odd index to create an even tensor. Sorting these dummy indices yields the appropriate global phase associated with the order of the odd tensors.
Clearly, not every locally ordered choice of arrow is consistent with a global ordering; for example, if the arrows form a cycle, then it is not a DAG. However, we can reverse the arrows, which is equivalent to placing a parity tensor on the reversed bond (Fig. 3b). If we do so to obtain a DAG of fermionic tensors (with parity tensors absorbed), this may be re-expressed in terms of the globally ordered contraction operation $\mathcal{F}$.
@@ IV Implementation
We have implemented the above fermionic tensor network algebra in an Abelian block-sparse and fermionic array library symmray [43], which acts as a new backend for quimb [9]. From a function perspective, the main changes are that we must redefine unary and binary tensor operations, such as transposition, fusion and contraction, to incorporate the new bond parity contributions. In the case of the globally ordered formalism, we also introduce additional bookkeeping to track the global position of the tensors in the total contraction $\mathcal{F}$. Otherwise, no significant changes are required, and our fermionic implementation can thus take advantage of the general tensor network infrastructure in quimb, including its ability to perform hyperoptimized exact and approximate contraction.
@@ V Examples
We consider the half-filled Hubbard model on the 3D diamond lattice. We approximate the ground-state by a PEPS tensor network with $\mathbb{U}(1)$ symmetry and optimize the ground-state using simple update imaginary time evolution [44] as implemented in quimb. The simulation is carried out at a fixed total particle number using $\mathbb{U}(1)$ symmetric tensors. The lattice is bipartite, so the auxiliary field quantum Monte Carlo (AFQMC) does not suffer from a sign problem and is exact up to statistical error. We use Dice [45] to compute the AFQMC ground-state energy as a benchmark.
@@
To evaluate the energy of the PEPS, we considered two strategies. The first is to use a direct approximate contraction $\langle\Psi|H|\Psi\rangle$ (we contract separately for each term in $H$) for the entire finite diamond graph using the hyperoptimized approximate contraction strategy in Sec. III.
We also compute the energy by a cluster approximation, as used, e.g., in Refs. [46, 47, 48, 13, 14]. In this case, for each local term $h_{t}$ in $H$, we define a cluster of tensors around the pair of sites on which $h_{t}$ acts out to a finite radius (graph distance) $r$. (We show a 2D example of this in Fig. 5). This approximation is motivated by the fact that contracting the environment of $h_{t}$ is the main cost in contracting the tensor network, but in a tensor network without loops that are appropriately gauged, this environment contraction just yields the identity. Thus, here, the environment tensors outside of the cluster are approximated by the simple update gauges on each of the dangling bonds. The contraction of the cluster can take advantage of the approximate contraction to reduce the cost further. In this work, we use $\chi=32$ for the approximate contraction of the cluster.
@@
We first show results for the $3\times 3\times 3$ and $4\times 4\times 4$ diamond lattices. We optimize a PEPS ground-state with $D=6$ and then compute the energy using approximate contraction (with the hyperoptimized path determined using default search parameters in quimb) with $\chi=32$ and $\chi=16$ for the two lattice sizes respectively, and compare against results from AFQMC in Fig. 6 and in Table 1. At the $D=6$ level, the PEPS energies are converged to within 1-2% of the exact AFQMC benchmark. From Fig. 6, we see that the PEPS energies are converging quite rapidly with $D$. The results are most accurate for larger $U$. This reflects the relative difficulty PEPS has in simulating more itinerant fermions at smaller $U$.
These smaller lattices also provide an opportunity for us to assess the cluster approximation for energy evaluation. In Fig. 7, we assess it as a function of radius $r$, comparing to the results from the full approximate contraction (with $\chi=32$ for the $3\times 3\times 3$ lattice and $\chi=16$ for the $4\times 4\times 4$ lattice). For the $3\times 3\times 3$ case, the cluster approximation achieves a relative error of less than 1% for $r\geq 2$; the errors are smaller for larger $U$, where the correlations are more local. For the $4\times 4\times 4$ structure, however, the smallest difference between the cluster approximation and the full approximate contraction actually occurs at $r=2$. Since the $4\times 4\times 4$ approximate contraction is itself not very accurate (as $\chi=16$ is quite small), this does not mean that the $r=2$ cluster gives the best result. Overall, however, the above suggests that the cluster approximation error with $r=2$ or $r=3$ is comparable to the accuracy of the ansatz itself for $D=6$, with the largest error appearing for $U=4$, the most itinerant system.
Figure 6: Ground-state energies for Hubbard model on $3\times 3\times 3$ (top row) and $4\times 4\times 4$ (bottom row) diamond lattice with $U=4,6,8$. The PEPS results, here computed using full approximate contraction, are compared with AFQMC.
@@
In Fig. 8, we show the PEPS ground-state energies with bond dimension up to $D=6$ on the $5\times 5\times 5$ lattice using the $r=3$ cluster approximation. We see a rapid decrease of the energy with $D$, similar to in the smaller lattices. The relative error compared to AFQMC is seen to be similar to the smaller lattice sizes, with the exception of $U=4$, where we have a slightly larger error of $\sim 3\%$.
Figure 8: Ground-state energies of Hubbard model on $5\times 5\times 5$ diamond lattice with $U=4,6,8$. The PEPS results, here computed with the $r=3$ cluster approximation, are compared with AFQMC.
@@
Table 1: Ground-state energies of the Hubbard model on diamond lattices. The first entry in the SU column is computed from approximate contraction of the whole lattice and the second from the cluster approximation.
Figure 9: Fermionic PEPS applied to the Fermi-Hubbard model on random 3-regular graphs with $U=8$ at half-filling. All lines show mean behavior across 20 instances with shaded areas denoting error on the mean. (a) Energy per site compared to exact diagonalization and DMRG as a function of number of sites, $N$, for varying PEPS dimension, $D$, and MPS bond dimension, $\chi$. (b) Error in energy compared to exact diagonalization. (c) Example PEPS geometry for a single instance at $N=64$.
@@ V.2 Random regular graphs
We next move on to the more unusual geometry of random 3-regular graphs, which are neither bipartite nor short-range and thus quite challenging for existing methods. We stick with the Hubbard model given by Eq. 18 in the strongly interacting regime of $U=8$ at half filling. Given the complex geometry, we again choose to evaluate the energy using the cluster approximation, which is generally non-variational but provides a useful overview of trends. We do know as the number of sites $N\rightarrow\infty$ that this approximation should approach exact however, given the graph loops become infinitely long, and the local geometry then becomes tree-like. We take 20 instances for each graph size $N$ and optimize fermionic PEPS with a variety of bond dimensions, $D$, using the simple update. The energy is then evaluated using the simple update gauges at cluster radius $r=0$. Up to $N=16$ we can compare to exact diagonalization, and beyond that we also compare to DMRG for a range of maximum bond dimensions $\chi$. We compute the DMRG results using the block2 package[49, 50], without spin adaptation. We converge each bond dimension $\chi$, then use that state as the initial guess for next, doubled $\chi$. At $N{=}64$ and $\chi{=}4096$ this requires multiple days of compute on 16 cores, compared to about ${\sim}15$ minutes for the $D{=}16$ PEPS computation.

### v3

@@ II.1 Standard tensor networks and symmetries
We will consider only symmetric tensors with well-defined total irrep. To define how the total irrep is obtained from the irreps of the indices, we label each index of the tensor with a plus (+, “outgoing”) or minus (-, “incoming”). Then, for example, if for $A_{ijkl}$ the labels of ijklijkl are ++{-}{+}{-}{+}, we have
$A_{ijkl}=0~~\mathrm{if}~~-s(i)+s(j)-s(k)+s(l)\neq s(A)$ (4)
where $s(A)$ is the total irrep of $A$. Similarly if ijklijkl are labelled ++++{+}{+}{+}{+}, the entries of $A_{ijkl}$ vanish if $s(i)+s(j)+s(k)+s(l)\neq s(A)$, and so on. Note that for finite Abelian groups (e.g., $\mathbb{Z}_{2}$), addition and subtraction is modulo $H$.
@@
The unordered contraction in Eq. 2 defines the value of $\mathcal{T}$. For symmetric tensors, the +/- labels of the contracted indices must be compatible for the contraction to be valid, that is, the outgoing leg of one tensor can only be contracted with the ingoing leg of another (in Eq. 1, either $k$ is labelled + on $A$ and - on $B$, or vice versa), and any such compatible choice of +/- labels is a valid symmetric tensor network $\mathcal{T}$. However, the cost to compute $\mathcal{T}$ further depends on the order of the tensor contractions. If approximate contraction is used, both cost and accuracy also depend on when and which shared bonds are compressed. The order of tensor contraction in the exact case defines a contraction tree (see Fig. 1(a)) where the vertical axis illustrates the time-ordering of the pairwise contractions. In the approximate contraction case, Fig. 1(b), the contraction tree contains additional lines between branches indicating the compression of bonds between tensors.
@@ II.2 Fermionic tensors and contractions
A fermionic tensor network is formed from fermionic tensors. We thus first discuss the properties of fermionic tensors. In our implementation, these are stored as Abelian arrays of numbers with additional metadata for parity and other symmetries, and the internal index order. Parity is a $\mathbb{Z}_{2}$ symmetry with irreps $\sigma\in\{0,1\}$; for index $i$, we write $p_{i}$ as the parity of $i$. We will assume the fermionic tensors have well-defined total parity 0 (even) or $1$ (odd). Fermionic tensors may have additional symmetries, which may or may not be independent of parity. A common symmetry is $\mathbb{U}(1)$ symmetry, associated with fermion number or projected spin number. In this case, odd integer irreps of $\mathbb{U}(1)$ have parity $1$, while even integer irreps of $\mathbb{U}(1)$ have parity 0.
@@ II.3 Fermionic tensor networks
We note that bosonic or spin degrees of freedom can be easily mixed into fermionic tensor networks by treating them as entirely even parity legs.
@@ II.4 Globally ordered formalism
where we store the additional data on the initialization order in the form of an integer position for each fermionic tensor (the superscripts in Eq. 14). $\mathcal{T}$ is then obtained by contracting the network in sequence from left to right. Transposition of $\mathcal{T}$ (e.g., when taking the hermitian conjugate to go from $|\Psi\rangle$ to $\langle\Psi|$) involves reversing the order of the tensors appearing in Eq. 14, as well and reversing the indices of all the tensors.
@@
In this DAG picture, we can also visualize what happens when one changes the global ordering. When moving tensors past each other to construct a new global order, one needs to construct a new DAG corresponding to the new order (see sequence in Fig. 3a). In general, this requires the reversal of multiple arrows in the original DAG, and these reversals correspond to the bond parity tensors that are discussed in Sec. II.2.
@@ II.5 Locally ordered formalism
From here, it is a short leap to the locally ordered formalism. It is clear from the above that it is not necessary to limit ourselves to DAGs, but one can use other arrow graphs, where the arrows only define the local ordering of two tensors in a given contraction. In this case, the value of the fermionic tensor network is given by
@@
where $\mathcal{G}$ denotes the fermionic contraction following the additional data where every contracted bond has an arrow, and there is no longer any global ordering of the tensors. The computation of the bond parity is needed only on the bond between two tensors when they are contracted: if they are contracted in the direction of the arrow that bonds them, there is no parity, while if they are contracted in the opposite order, we compute the bond parity (or equivalently, first reverse the arrow). The above assumes that we are working with even parity tensors, so there is no global phase. However, if there are odd total parity tensors in $\mathcal{T}$, we also need to track the phase associated with the ordering of the odd parity tensors. It is convenient to do so by implementing odd parity tensors using an additional dummy (size 1) odd index to create an even parity tensor. The dummy indices are then propagated to the final tensor that is the result of the full contraction of the tensor network, and sorting the remaining dummy indices according to a chosen global ordering convention yields the phase defined by the odd parity tensors (according to the same convention).
Clearly, not every locally ordered choice of arrows is consistent with a global ordering; for example, if the arrows form a cycle, then it is not a DAG. However, we can reverse the arrows, which is equivalent to placing a parity tensor on the reversed bond (Fig. 3b). If we do so to obtain a DAG of fermionic tensors (with parity tensors absorbed), this may be re-expressed in terms of the globally ordered contraction operation $\mathcal{F}$.
@@ IV Implementation
We have implemented the above fermionic tensor network algebra in an Abelian block-sparse and fermionic array library symmray [43], which acts as a new backend for quimb [9]. No significant changes are required, and our fermionic implementation can thus take advantage of the general tensor network infrastructure in quimb, including its ability to perform hyperoptimized exact and approximate contraction. Below, we discuss some more specific implementation details.
IV.1 Tensor operations and index fusion
From a function perspective, the main changes are that we must redefine unary and binary tensor operations, such as transposition and contraction, to incorporate the new bond parity contributions. If using a global ordering throughout the calculation, we introduce additional bookkeeping to track the global position of the tensors in the total contraction $\mathcal{F}$.
In addition, we also implement fusion (combining multiple indices into one, either for efficient contraction or linear algebra decompositions), which involves multiple parity contributions and is worth a small discussion. The first arises from the fermionic transposition to make the indices adjacent. Second, all arrows in the fusion set are required to point in the same direction (as the first arrow), which may require flipping arrows. Third, for a fused bond with an inward arrow, we reverse the ordering of the fermionic modes so we obtain the same parity contributions as if the modes were contracted one by one, i.e. $\sum_{ijk}A_{m(ijk)}B_{(ijk)n}=\sum_{i}\sum_{j}\sum_{k}{A}_{m(ijk)}\tilde{B}_{(kji)n}$, where $(\ldots)$ denotes the fused index, and $\tilde{B}$ indicates that the parity contributions of $(ijk)\to(kji)$ have been accumulated into tensor $B$.
IV.2 Tensor network initialization
When initializing an arbitrary tensor network, the direction of the arrows can be chosen completely arbitrarily (locally ordered formalism) or they can be inferred from a given global ordering of the tensors (globally ordered formalism). In practice, by default, we initialize the arrows so that they point from later to earlier sites, which is consistent with a global ordering, and a DAG. In principle, there is an additional choice of the incoming/outcoming symmetry label of each tensor leg (which must also be done consistently), but we can choose this labelling to align with the arrow directions (i.e. the arrow points from ++ to -), and this is done by default.
Figure 4: (a) Norm tensor network for $|\Psi\rangle=\mathcal{F}(A^{[0]}B^{[1]}C^{[2]})$ as $\langle\Psi|\Psi\rangle=\mathcal{F}(C^{{\dagger}[0]}B^{{\dagger}[1]}A^{{\dagger}[2]}A^{[3]}B^{[4]}C^{[5]})$. (b) Norm for the partial ‘cluster’ formed by taking only the central site $\mathcal{F}(B^{{\dagger}[0]}B^{[1]})$, requiring some change of arrow directions.
As an example of this initialization, we consider the case of setting up the tensor network corresponding to the norm $\langle\Psi|\Psi\rangle$. For a $|\Psi\rangle$ given by $\mathcal{F}(A^{[0]}B^{[1]}C^{[2]})$, we can obtain $\langle\Psi|$ by transposition, giving the overall globally ordered tensor network $\mathcal{F}(C^{{\dagger}[0]}B^{{\dagger}[1]}A^{{\dagger}[2]}A^{[3]}B^{[4]}C^{[5]})$. Note that the tensors in the bra all occur before those in the ket, and inferring the arrow directions of the DAG from the global ordering yields the picture shown in Fig. 4(a), where the bra arrows are all the reverse of those in the ket.
As a second example, in the cluster algorithm implemented later in Sec. V, we extract part of the tensor network (a cluster) in the bra and contract it with the same cluster in the ket; e.g. consider extracting only the $B$ tensors from the tensor network expression above. The norm then becomes the tensor network $\mathcal{F}(B^{{\dagger}[0]}B^{[1]})$, corresponding to the DAG in Fig. 4(b). Note that the arrows joining the $B$ tensors in Fig. 4(b), as inferred from the DAG, are different to the arrows joining the $B$ tensors in Fig. 4(a), as some are reversed. This is equivalent to implementing the bond parity tensor coming from “partial trace” operations discussed in Ref. [32].
@@ V Examples
We consider the half-filled Hubbard model on the 3D diamond lattice. We approximate the ground-state by a PEPS tensor network with $\mathbb{U}(1)$ symmetry and optimize the ground-state using simple update (SU) imaginary time evolution [44] as implemented in quimb. The simulation is carried out at a fixed total particle number using $\mathbb{U}(1)$ symmetric tensors. The lattice is bipartite, so the auxiliary field quantum Monte Carlo (AFQMC) does not suffer from a sign problem and is exact up to statistical error. We use Dice [45] to compute the AFQMC ground-state energy as a benchmark.
@@
To evaluate the energy of the PEPS, we considered two strategies. The first is to use a direct approximate contraction $\langle\Psi|H|\Psi\rangle$ (we contract separately for each term in $H$) for the entire finite diamond graph using the hyperoptimized approximate contraction strategy in Sec. III. The second is to use a cluster approximation, detailed below.
Figure 6: Ground-state energies for the Hubbard model on the $3\times 3\times 3$ (top row) and $4\times 4\times 4$ (bottom row) diamond lattice with $U=4,6,8$. The PEPS results, which are optimized with SU, and with the energy computed using full approximate contraction, are compared with AFQMC.
We first show results for the $3\times 3\times 3$ and $4\times 4\times 4$ diamond lattices using energies evaluated by hyperoptimized approximate contraction. We optimize a PEPS ground-state with $D=6$ and for the approximate contraction hyperoptimized path we use the default search parameters in quimb with $\chi=32$ and $\chi=16$ for the two lattice sizes respectively. We compare against exact results from AFQMC in Fig. 6 and in Table 1. At the $D=6$ level, the PEPS energies are converged to within 1-2% of the exact AFQMC benchmark. From Fig. 6, we see that the PEPS energies are converging quite rapidly with $D$. The results are most accurate for larger $U$. This reflects the relative difficulty PEPS has in simulating more itinerant fermions at smaller $U$.
@@
We next compute the energy by a cluster approximation, as used, e.g., in Refs. [46, 47, 48, 13, 14]. In this case, for each local term $h_{t}$ in $H$, we define a cluster of tensors around the pair of sites on which $h_{t}$ acts out to a finite radius (graph distance) $r$. (We show a 2D example of this in Fig. 7). This approximation is motivated by the fact that contracting the environment of $h_{t}$ is the main cost in contracting the tensor network, but in a tensor network without loops that are appropriately gauged, this environment contraction just yields the identity. Thus, here, the environment tensors outside of the cluster are approximated by the simple update gauges on each of the dangling bonds. The contraction of the cluster can take advantage of the approximate contraction to reduce the cost further. In this work, we use $\chi=32$ for the approximate contraction of the cluster.
The smaller lattices provide an opportunity for us to assess the cluster approximation for energy evaluation. In Fig. 8, we assess it as a function of radius $r$, comparing to the results from the full approximate contraction (with $\chi=32$ for the $3\times 3\times 3$ lattice and $\chi=16$ for the $4\times 4\times 4$ lattice). For the $3\times 3\times 3$ case, the cluster approximation achieves a relative error of less than 1% for $r\geq 2$; the errors are smaller for larger $U$, where the correlations are more local. For the $4\times 4\times 4$ structure, however, the smallest difference between the cluster approximation and the full approximate contraction actually occurs at $r=2$. Since the $4\times 4\times 4$ approximate contraction is itself not very accurate (as $\chi=16$ is quite small), this does not mean that the $r=2$ cluster gives the best result. Overall, however, the above suggests that the cluster approximation error with $r=2$ or $r=3$ is comparable to the accuracy of the ansatz itself for $D=6$, with the largest error appearing for $U=4$, the most itinerant system.
@@
In Fig. 9 and Table 1, we show the SU optimized PEPS ground-state energies with bond dimension up to $D=6$ on the $5\times 5\times 5$ lattice where the final energy is evaluated using the $r=3$ cluster approximation. We see a rapid decrease of the energy with $D$, similar to in the smaller lattices. The relative error compared to AFQMC is seen to be similar to the smaller lattice sizes, with the exception of $U=4$, where we have a slightly larger error of $\sim 3\%$.
Figure 9: Ground-state energies of the Hubbard model on the $5\times 5\times 5$ diamond lattice with $U=4,6,8$. The PEPS results, which are optimized with SU, with the final energy computed with the $r=3$ cluster approximation, are compared with AFQMC.
@@
Table 1: Ground-state energies of the Hubbard model on diamond lattices. The first entry in the SU column is computed from approximate contraction of the whole lattice and the second from the cluster approximation with $r=3$.
Figure 10: Fermionic PEPS applied to the Fermi-Hubbard model on random 3-regular graphs with $U=8$ at half-filling. The PEPS is optimized with SU, with energy computed using the $r=0$ cluster approximation. All lines show mean behavior across 20 instances with shaded areas denoting error on the mean. (a) Energy per site compared to exact diagonalization and DMRG as a function of number of sites, NN, for varying PEPS dimension, $D$, and MPS bond dimension, $\chi$. (b) Error in energy compared to exact diagonalization. (c) Example PEPS geometry for a single instance at $N=64$.
@@ V.2 Random regular graphs
We next move on to the more unusual geometry of random 3-regular graphs, which are neither bipartite nor short-range and thus quite challenging for existing methods. We stick with the Hubbard model given by Eq. 18 in the strongly interacting regime of $U=8$ at half filling. Given the complex geometry, we again choose to evaluate the energy using the cluster approximation, which is generally non-variational but provides a useful overview of trends. We do know as the number of sites $N\rightarrow\infty$ that this approximation should become exact however, given the graph loops become infinitely long, and the local geometry then becomes tree-like. We take 20 instances for each graph size $N$ and optimize fermionic PEPS with a variety of bond dimensions, $D$, using the simple update. The energy is then evaluated using the simple update gauges at cluster radius $r=0$. Up to $N=16$ we can compare to exact diagonalization, and beyond that we also compare to DMRG for a range of maximum bond dimensions $\chi$. We compute the DMRG results using the block2 package[49, 50], without spin adaptation. We converge each bond dimension $\chi$, then use that state as the initial guess for the next, doubled $\chi$. At N=64N{=}64 and $\chi{=}4096$ this requires multiple days of compute on 16 cores, compared to about ${\sim}15$ minutes for the D=16D{=}16 PEPS computation.